OpenStudy (shubhamsrg):

integral [ log( sqrt(1+x) + sqrt(1-x) ) dx] from x=0 to 1

OpenStudy (lgbasallote):

\[\huge \int \limits_0^1 \log (\sqrt{1 + x} + \sqrt{1- x})\] ???

OpenStudy (anonymous):

yep i think so

OpenStudy (shubhamsrg):

yep.. and a dx there,,ofcorse..

OpenStudy (anonymous):

ah.. leave the typos.. concentrate on question. lol

hartnn (hartnn):

simple substitution of x=cos 2t

OpenStudy (shubhamsrg):

then our dx = -2sin 2t dt.. dont you think that will create a problem ? altough the sq. roots will be eliminated..

OpenStudy (shubhamsrg):

how'd you get that ? what i got was inside the integral (something) + log( cos t + sin t ) sin 2t dt are you missing the log there buddy ?

OpenStudy (anonymous):

is that \(\ln\)

OpenStudy (shubhamsrg):

yep..ln..

hartnn (hartnn):

i need to reconsider

OpenStudy (anonymous):

man this is interesting

OpenStudy (anonymous):

let\[t=\ln(\sqrt{1+x}+\sqrt{1-x})\]

hartnn (hartnn):

is there a standard formula for: \[\int\limits_{}^{}\ln(1+\sqrt{1-x^2})\] if yes,then this is so simple!

OpenStudy (anonymous):

i think there is

hartnn (hartnn):

\[=(1/2)\int\limits \limits_0^1 \log (\sqrt{1 + x} + \sqrt{1- x})^2 dx\] \[=(1/2)\int\limits \limits_0^1 \log (2+2\sqrt{1-x^2})^2 dx\] \[=(1/2)\int\limits \limits_0^1 \ln 2+\ln (1+\sqrt{1-x^2})^2 dx\]

hartnn (hartnn):

\[=(1/2)\int\limits\limits\limits \limits_0^1 \ln 2+\int\limits\limits\limits \limits_0^1\ln (1+\sqrt{1-x^2}) dx\]

OpenStudy (anonymous):

\[\ln(1+\sqrt{1-x^2})=\text{sech}^{-1}x+\ln x\]this is hard to get o.O

hartnn (hartnn):

lol,i did same exact thing :P

hartnn (hartnn):

i guess,thats integration by parts........???

hartnn (hartnn):

taking u=ln(....),v=1

OpenStudy (shubhamsrg):

bingo,,got it by parts! @hartnn thanks a lot!

hartnn (hartnn):

belated welcome :)